Please help me answer the below one question (SH - highlighted in yellow. The answer is not 38,000:

Antuan Company set the following standard costs for one unit of its product. Direct materials (5.0 Ibs. $6.00 per Ib.) $30.00 Direct labor (1.9 hrs. $12.00 per hr.) Overhead (1.9 hrs. $18.50 per hr.) Total standard cost 22.80 35.15 $87.95 The predetermined overhead rate ($18.50 per direct labor hour) is based on an expected volume of 75% of the factorys capacity of 20,000 units per month. Following are the companys budgeted overhead costs per month at the 75% capaci level Overhead Budget (75% Capacity) Variable overhead costs Indirect materials Indirect labor Power Repairs and maintenance s 15,000 75,000 15,000 30,000 Total variable overhead costs - $135,000 Fixed overhead costs Depreciation-Building Depreciation-Machinery Taxes and insurance Supervision Total fixed overhead costs 25,000 70,000 17,000 280,250 392,250 $527,250 Total overhead costs The company incurred the following actual costs when it operated at 75% of capacity in October Direct materials (76, 500 Ibs. $6.20 per S 474, 300 Direct labor (28,000 hrs. $12.30 per hr.) Overhead costs 344, 400 Indirect materials Indirect labor Power Repairs and maintenance Depreciation-Building Depreciation-Machinery Tazes and insurance Supervision S 41,300 176, 050 17,250 34, 500 25,000 94, 500 15,300 280,250 684,150 $1,502,850 Total Co3ts
Answer #1
SH or Standard hours = 15000*1.9 = 28500
Standard cost = 28500*12= 342000
Direct labor efficiency variance =342000-336000 = $6000 favorable